Cindy is painting a rectangular prism. The width is 6 inches ,the height is 5 inches and the length is 7inches . What is the surface area of the rrectangular prism?
A)107 square inches
B)214 square inches
C)210 square inches
D)154 square inches

Answer by jim_thompson5910(35256)

SA = 2*(L*W + W*H + L*H)

SA = 2*(7*6 + 6*5 + 7*5)

SA = 2*(42 + 30 + 35)

SA = 2*(107)

SA = 214 square inches